# ArkTS-CodeSearch: A Open-Source ArkTS Dataset for Code Retrieval

### v1

@@ 5.2. Fine-tuning on ArkTS Dataset
Figure 7. Model size versus MRR before and after fine-tuning on the ArkTS dataset.
The Qwen/Qwen3-Embedding-0.6B model nonetheless shows a substantial absolute improvement, increasing its MRR from 0.6776 to 0.7745. This demonstrates that large-scale pretrained models remain highly capable, but still benefit significantly from ArkTS-specific supervision.
The multilingual BAAI/bge-m3 model ranks third after fine-tuning, achieving an MRR of 0.7467 and Recall@5 of 0.8504, reflecting strong generalization ability across languages and tasks.
Importantly, smaller models such as BAAI/bge-base-zh-v1.5 exhibit pronounced performance gains after fine-tuning, narrowing the gap with substantially larger models. This observation suggests that, for ArkTS code retrieval, domain adaptation and task-specific training objectives are more critical than model scale alone.
In contrast, English-oriented or general-purpose sentence embedding models, including BAAI/bge-base-en-v1.5 and intfloat/e5-base-v2, remain less competitive even after fine-tuning, highlighting the combined impact of domain mismatch and language mismatch.
Overall, these results demonstrate that supervised contrastive fine-tuning on ArkTS-specific data is highly effective and can significantly alter the relative performance of embedding models. The optimized models consistently outperform their zero-shot baselines, underscoring the importance of domain- and language-aware datasets for semantic code search.
Models MRR NDCG RECALL@1 RECALL@5
google/embeddinggemma-300m 0.7788 0.8034 0.7142 0.8769
Qwen/Qwen3-Embedding-0.6B 0.7745 0.7971 0.7146 0.8643
BAAI/bge-m3 0.7467 0.7729 0.6750 0.8504
BAAI/bge-base-zh-v1.5 0.6748 0.7055 0.5961 0.7972
BAAI/bge-base-en-v1.5 0.4866 0.5144 0.4195 0.5977
intfloat/e5-base-v2 0.4855 0.5135 0.4158 0.5973
Table 2. Leaderboard for finetuned opensorce models(ARKTS-TRAINING)
@@
6. Discussion
The experimental results and dataset analysis provide several insights into ArkTS code semantic search and code representation learning.
First, the dataset statistics reveal substantial diversity in both code structure and documentation practices. While most ArkTS functions are relatively short and lightly documented, the presence of long-tailed distributions in AST size, function length, and docstring length indicates that a non-negligible portion of the codebase contains complex logic and detailed explanations. The observed secondary peak in docstring length further suggests heterogeneous documentation patterns, where core modules or publicly exposed interfaces tend to receive more extensive comments. These characteristics pose challenges for semantic code search models, which must handle both concise and richly documented code.
Second, the zero-shot evaluation highlights the limitations of directly applying off-the-shelf embedding models to ArkTS code retrieval. Although large-scale and multilingual models such as Qwen3 and embeddinggemma demonstrate relatively strong performance, general-purpose sentence embedding models perform substantially worse, reflecting a clear domain mismatch. In particular, among models with similar parameter scales, Chinese-oriented embedding models outperform their English counterparts, likely due to the prevalence of Chinese-language docstrings in the ArkTS ecosystem. This observation underscores the importance of language alignment between training data and downstream tasks.
Third, fine-tuning on ArkTS-specific data consistently yields significant performance gains across all models and fundamentally alters their relative rankings. Mid-sized models, when properly adapted, can surpass larger pretrained models, indicating that domain-specific supervision plays a more critical role than model scale alone. These results demonstrate that supervised contrastive learning with high-quality, task-aligned data is an effective strategy for ArkTS code retrieval.
Overall, our findings suggest that the combination of a well-curated dataset, language-aware modeling, and task-specific fine-tuning is essential for advancing code intelligence research on emerging programming languages such as ArkTS.
@@ 7. Conclusion and Future Work
In this work, we present the first large-scale, publicly available dataset and benchmark for ArkTS code semantic search. By collecting ArkTS repositories from both GitHub and Gitee and leveraging AST-based parsing with tree-sitter-arkts, we construct a high-quality dataset of aligned docstring–code pairs with rich statistical characterization. We further provide a comprehensive evaluation of existing open-source embedding models and demonstrate that domain-specific fine-tuning substantially improves retrieval performance.
Our experimental results show that, while large pretrained embedding models perform reasonably well in a zero-shot setting, supervised contrastive fine-tuning on ArkTS-specific data is crucial for achieving strong performance. Moreover, the observed performance gains of mid-sized and language-aligned models indicate that effective domain adaptation can outweigh model scale, particularly for emerging programming languages with unique linguistic and ecosystem characteristics.
There are several promising directions for future work. First, the dataset can be extended to support additional tasks such as commit-based code retrieval, code change understanding, and software evolution analysis by incorporating commit messages and diffs. Second, future studies may explore structure-aware models that explicitly leverage AST or graph-based representations to further improve semantic alignment between code and natural language. Finally, expanding the dataset to include more diverse OpenHarmony projects and multilingual annotations could facilitate broader research on cross-language and cross-platform code intelligence.
We hope that this dataset and benchmark will serve as a foundation for future research on ArkTS and contribute to the development of more effective code intelligence techniques for the OpenHarmony ecosystem.

### v2

@@ 5.2. Fine-Tuning
Figure 7. Model size versus MRR before and after fine-tuning on the ArkTS dataset. A small horizontal jitter is applied for clarity of visualization.
Fine-tuning substantially improves retrieval performance across all embedding models compared to zero-shot baselines, demonstrating the effectiveness of task-specific supervision for aligning natural-language docstrings with ArkTS code. Evaluation results are summarized in Tables 1 and 2, and a well-performing fine-tuned model trained on the ArkTS training set is publicly released on Hugging Face. [^8] Available at https://huggingface.co/hreyulog/embedinggemma_arkts The best-performing model is available upon reasonable request.
The choice of training data has a pronounced impact on retrieval quality. Models adapted to ArkTS-specific supervision achieve consistently higher scores than those trained only on the TypeScript corpus, across all metrics including MRR, NDCG, Recall@1, and Recall@5. This highlights the critical role of domain-aligned fine-tuning in semantic code retrieval.
The proposed TS→ArkTS training strategy improves performance over single-stage fine-tuning in most cases, especially for models with larger parameter sizes. The two-stage procedure benefits from the broader semantic coverage of TypeScript data followed by ArkTS-specific adaptation, leading to strong gains in early-recall metrics. In contrast, models with approximately 100M parameters do not consistently benefit from this strategy and may experience performance degradation, potentially due to limited model capacity under successive fine-tuning stages.
Among all evaluated models, google/embeddinggemma-300m achieves the best performance under all three training configurations, reaching an MRR of 0.6259 with TS-TRAINING and 0.8013 with TS→ARKTS-TRAINING. Notably, this mid-sized model outperforms larger alternatives such as Qwen/Qwen3-Embedding-0.6B, indicating that effective domain adaptation can outweigh advantages from increased model scale. Nevertheless, Qwen/Qwen3-Embedding-0.6B exhibits substantial absolute improvements after ArkTS-specific fine-tuning, demonstrating that large pretrained models also benefit significantly from targeted supervision.
Multilingual and Chinese-oriented models, including BAAI/bge-m3 and BAAI/bge-base-zh-v1.5, show strong and stable gains across all training settings, particularly in Recall@1, reflecting improved precision for top-ranked results. In contrast, English-centric sentence embedding models such as BAAI/bge-base-en-v1.5 and intfloat/e5-base-v2 remain less competitive. We attribute this difference to the linguistic characteristics of the ArkTS corpus: a substantial portion of code comments and docstrings are written in Chinese. As a result, models pretrained with stronger Chinese language representations are better aligned with the natural-language queries and documentation present in the dataset, leading to improved semantic matching.
Overall, these results demonstrate that training data composition is as critical as model architecture for semantic code search. Supervised contrastive fine-tuning on ArkTS-specific data substantially reshapes model rankings, and the proposed TS→ARKTS two-stage strategy provides a practical and effective approach for adapting general-purpose embedding models to specialized programming languages.
@@ 6. Conclusion and future work
Our experiments and dataset analysis provide several key insights into ArkTS code semantic search and representation learning. The ArkTS dataset exhibits substantial diversity in both code structure and documentation practices: while most functions are short and lightly documented, long-tailed distributions in AST size, function length, and docstring length indicate that a non-negligible subset of the codebase involves complex logic and detailed explanations. The prevalence of Chinese-language docstrings further highlights the importance of language-aware modeling, as Chinese-oriented embeddings consistently outperform English-oriented counterparts in retrieval tasks.
[…]
Looking forward, several directions are promising. Incorporating commit messages, diffs, and additional OpenHarmony projects could extend the dataset to support commit-based code retrieval, code change understanding, and software evolution analysis. Structure-aware models that explicitly leverage AST or graph-based representations may further enhance semantic alignment between code and natural language. Expanding multilingual annotations would also facilitate broader cross-language and cross-platform code intelligence research. We hope this dataset and benchmark will serve as a foundation for future research on ArkTS and contribute to the development of more effective code intelligence techniques within the OpenHarmony ecosystem.